OpenStudy (anonymous):

helppppppppppp!! Two times Owen’s age plus 4 times Charlotte’s age equals 42. Charlotte’s age is also 3 times Owen’s age. How old is Charlotte?

OpenStudy (ivettef365):

you have Owen = W Charlotte = C 2W +4C = 42 C = 3W

OpenStudy (ivettef365):

so you can substitute the C on first equation for 3W and you have 2W + 4(3W) = 42 now solve for W, then find what is C

OpenStudy (anonymous):

Let x : Owen's age and y: Charlotte's age so 3x+4y=42 and y=3x ==> 3x+4*3x=42 ==> 3x+12x =42 ==> x=42/15 ==> y= (3*42)/15

OpenStudy (anonymous):

how'd you get 42

OpenStudy (anonymous):

Two times Owen’s age plus 4 times Charlotte’s age equals 42

OpenStudy (ivettef365):

so did you get what is Charlotte's age ?

OpenStudy (anonymous):

is it 6

OpenStudy (anonymous):

wait no

OpenStudy (anonymous):

4(3) is 12, then what do you do

OpenStudy (ivettef365):

no, lets solve the problem, 2W + 12W = 42 14W = 42 2 = 3 so, Owen's age is 3 years old now you know that Charlotte is 3 times Owens age so what is Charlottes age

OpenStudy (ivettef365):

I meant W = 3

OpenStudy (anonymous):

its wrong it was 9

OpenStudy (ivettef365):

what is Charlottes age ?

OpenStudy (ivettef365):

correct

OpenStudy (ivettef365):

since Owen's age is 3 Charlotte's age is 3 times Owen then Charlotte's age is 3 x 3 = 9

OpenStudy (anonymous):

thanks

OpenStudy (ivettef365):

yw :)